# Communication-Efficient Distributed Deep Learning via Federated Dynamic Averaging

### v1

@@ 3. Federated Dynamic Averaging
(6) $\;H(S(t))\leq\Theta\;\Rightarrow\;\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)\leq\Theta\quad\forall\;t$
Conceptually, if we define $S_{k}(t)=\left(\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}\,,\,\mathbf{u}_{t}^{(k)}\right)\in\mathbb{R}\times\mathbb{R}^{d}$ and $H(v,x)=v-\left\|x\right\|_{2}^{2}$, then, it immediately follows from (3), (5) that $H\left(S\left(t\right)\right)=\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$ and property (6) is satisfied. Direct monitoring of the RI is hindered by the high dimensionality of $S_{k}(t)$, with $d$ potentially ranging from millions to many billions, values commonly found in DNNs. To mitigate this communication burden, we must apply dimensionality reduction to the local drifts $\mathbf{u}_{t}^{(k)}$ and identify an appropriate function $H$, accordingly. However, this reduction in information within local states causes RI monitoring to become approximate. Upcoming sections will detail two strategies–termed ”linear”, and ”sketch”–each offering a different balance between communication efficiency and approximation accuracy.
Algorithm. The proposed FDA algorithm is formalized in Algorithm 1. To streamline notation, $t$ also denotes each training round, which unfolds into three main parts:
  1. (1)
    Broadcast: Model $\mathbf{w}_{t}$ is distributed to all workers (Line 3).
  2. (2)
    Local Training: As long as we can guarantee the RI, workers continue training concurrently (Lines 4-9). Specifically, workers perform in-parallel updates to their local models $\mathbf{w}^{(k)}_{t}$ using mini-batch SGD, or some other optimization algorithm, by processing their local data (Lines 5-7). After each update, workers compute and broadcast their local states $S_{k}(t)$ (Line 8). Then, the global state $S(t)$ is calculated via (5), and using the function $H$, an estimate of the model variance is derived and compared with the threshold $\Theta$ (Line 9). Depending on this comparison, the local training stage either repeats or terminates.
  3. (3)
    Model Averaging: This stage is triggered when we can no longer guarantee the RI, i.e., $H(S(t))>\Theta$. At this point, all local models are communicated and then averaged to form $\mathbf{w}_{t+1}$, that is, the next round’s global model (Lines 10-12).
@@
1:Require: $S_{k}(t)$: The local state $S_{k}(t)\in\mathbb{R}\times\mathbb{R}^{p}$ with $p\ll d$
2:Require: $H(x)$: A function s.t. $H(S(t))\leq\Theta\Rightarrow\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)\leq\Theta$
3:Require: $K$: The number of workers indexed by $k$
4:Require: $\Theta$: The model variance threshold
5:Require: $b$: The local mini-batch size
6:
7:Initialize $\mathbf{w}_{1}\in\mathbb{R}^{d}$
8:for each round $t=1,2,\dots$ do
9:     communicate $\mathbf{w}_{t}$ to all workers
10:     repeat
11:          for each worker $k=1,\dots,K$ in parallel do
12:               $\mathcal{B}_{k}\leftarrow$ (sample a batch of size $b$ from $\mathcal{D}_{k}$)
13:               $\mathbf{w}^{(k)}_{t}\leftarrow\mathbf{w}^{(k)}_{t}-\eta\nabla\ell(\mathbf{w}^{(k)}_{t};\mathcal{B}_{k})$triangleright In-place
14:          communicate $S_{k}(t)$
15:     until $H(S(t))>\Theta$triangleright As per (5)
16:     for each worker $k=1,\dots,K$ in parallel do
17:          communicate $\mathbf{w}^{(k)}_{t}$
18:     $\mathbf{w}_{t+1}\leftarrow\frac{1}{K}\sum_{k=1}^{K}\mathbf{w}^{(k)}_{t}$
19:
3.1. Linear Approximation
[…]
$S_{k}(t)=\left(\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}\,,\,\left\langle\xi\,,\,\mathbf{u}_{t}^{(k)}\right\rangle\right)\in\mathbb{R}\times\mathbb{R}\;\;,\;\;\left\|\xi\right\|_{2}=1$
$H\left(v\,,\,x\right)=v-x^{2}$
Then, the condition $H\left(S\left(t\right)\right)\leq\Theta$ implies $\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)\leq\Theta$, since:
$H\left(S\left(t\right)\right)$ =(5)${=}H\left(\frac{1}{K}\sum_{k=1}^{K}S_{k}(t)\right)=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left|\frac{1}{K}\sum_{k=1}^{K}\left\langle\xi\,,\,\mathbf{u}_{t}^{(k)}\right\rangle\right|^{2}$
$=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left|\left\langle\xi\,,\,\overline{\mathbf{u}}_{t}\right\rangle\right|^{2}\geq\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}=\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$
Furthermore, a random choice of $\xi$ is likely to perform poorly (terminate a round prematurely), as it is likely to be close to orthogonal to $\overline{\mathbf{u}}_{t}$. A good choice would be a vector correlated to $\overline{\mathbf{u}}_{t}$. A heuristic choice is to take $\overline{\mathbf{u}}_{t_{0}}$, i.e., the change vector right before the current round started. All workers can estimate this without communication cost as the difference of the models at the beginning of the current and previous rounds and use $\xi=\frac{w_{t_{0}}-w_{t_{-1}}}{\left\|w_{t_{0}}-w_{t_{-1}}\right\|_{2}}.$
3.2. Sketch Approximation
A better estimator for $\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$ can be obtained through the utilization of AMS sketches, as detailed in (Cormode and Garofalakis, 2005). An AMS sketch of a vector $\mathbf{v}\in\mathbb{R}^{d}$ is an $l\times m$ real matrix $\Xi$:
$\operatorname{sk}\left(\mathbf{v}\right)=\Xi=\begin{bmatrix}\xi_{1}&\xi_{2}&\dots&\xi_{l}\end{bmatrix}^{\top}\in\mathbb{R}^{l\times m}\;\;,\;\;l\cdot m\ll d$
@@ 3.2. Sketch Approximation
$\mathcal{M}_{2}\left(\Xi\right)=\mathrm{median}\left\{\left\|\xi_{i}\right\|_{2}^{2}\;,\;i=1,\dots,l\right\}$
@@
is an excellent estimator of $\left\|\mathbf{v}\right\|_{2}^{2}$, as with probability at least $\left(1-\delta\right)$, $\mathcal{M}_{2}\left(\operatorname{sk}\left(\mathbf{v}\right)\right)$ is within $\epsilon$-relative error of $\left\|\mathbf{v}\right\|_{2}^{2}$. We then define:
@@
$S_{k}(t)=\left(\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}\,,\,\operatorname{sk}\left(\mathbf{u}_{t}^{(k)}\right)\right)\in\mathbb{R}\times\mathbb{R}^{l\times m}$
$H\left(\mathbf{v}\,,\,\Xi\right)=\mathbf{v}-\frac{1}{1+\epsilon}\mathcal{M}_{2}\left(\Xi\right)$
Then, the condition $H\left(S\left(t\right)\right)\leq\Theta$ implies $\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)\leq\Theta$ with probability at least $\left(1-\delta\right)$. To prove this, we proceed as follows: we first observe $\mathcal{M}_{2}\left(\frac{1}{K}\sum_{k=1}^{K}\operatorname{sk}\left(\mathbf{u}_{t}^{(k)}\right)\right)=\mathcal{M}_{2}\left(\operatorname{sk}\left(\frac{1}{K}\sum_{k=1}^{K}\mathbf{u}_{t}^{(k)}\right)\right)=\mathcal{M}_{2}\left(\operatorname{sk}\left(\overline{\mathbf{u}}_{t}\right)\right)\leq(1+\epsilon)\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$ with probability at least $(1-\delta)$. Then:
@@
$H\left(S\left(t\right)\right)$ $=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\frac{1}{1+\epsilon}\mathcal{M}_{2}\left(\frac{1}{K}\sum_{k=1}^{K}\operatorname{sk}\left(\mathbf{u}_{t}^{(k)}\right)\right)$
@@
$\geq\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}=\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)\;\text{with prob. at least}\;(1-\delta)$

### v2

@@ 3. Federated Dynamic Averaging
(4) $\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)=\left(\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}^{(k)}_{t}\right\|_{2}^{2}\right)-\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$
Proof. Adding an offset ($-\overline{\mathbf{w}}_{t_{0}}$) to each $\mathbf{w}_{t}^{(k)}$ does not alter the variance, therefore:
$\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$ $=\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}-\overline{\mathbf{w}}_{t_{0}}\right)=\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{u}_{t}\right)=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}-\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$
$=\frac{1}{K}\sum_{k=1}^{K}\left(\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-2\left\langle\mathbf{u}_{t}^{(k)}\,,\,\overline{\mathbf{u}}_{t}\right\rangle+\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}\right)$
$=\left(\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}^{(k)}_{t}\right\|_{2}^{2}\right)-2\left(\frac{1}{K}\sum_{k=1}^{K}\left\langle\mathbf{u}_{t}^{(k)}\,,\,\overline{\mathbf{u}}_{t}\right\rangle\right)+\left(\frac{1}{K}\sum_{k=1}^{K}\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}\right)$
$=\left(\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}^{(k)}_{t}\right\|_{2}^{2}\right)-2\left\langle\left(\frac{1}{K}\sum_{k=1}^{K}\mathbf{u}_{t}^{(k)}\right)\,,\,\overline{\mathbf{u}}_{t}\right\rangle+\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$
$=\left(\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}^{(k)}_{t}\right\|_{2}^{2}\right)-2\left\langle\overline{\mathbf{u}}_{t}\,,\,\overline{\mathbf{u}}_{t}\right\rangle+\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$
$=\left(\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}^{(k)}_{t}\right\|_{2}^{2}\right)-2\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}+\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$
$=\left(\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}^{(k)}_{t}\right\|_{2}^{2}\right)-\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$
@@
1:Require: $K$: The number of workers indexed by $k$
2:Require: $\Theta$: The model variance threshold
3:Require: $b$: The local mini-batch size
4:
5:Initialize $\mathbf{w}_{0}^{(k)}=\overline{\mathbf{w}}_{0}\in\mathbb{R}^{d}$
6:for each step $t=1,2,\dots$ do
7:     for each worker $k=1,\dots,K$ in parallel do
8:          $\mathcal{B}_{t}^{(k)}\leftarrow$ (sample a batch of size $b$ from $\mathcal{D}_{k}$)
9:          $\mathbf{w}^{(k)}_{t}\leftarrow\textsc{Optimize}(\mathbf{w}_{t-1}^{(k)}\,,\,\mathcal{B}_{t}^{(k)})$
10:          Update $S_{t}^{(k)}$
11:          $\overline{\mathbf{S}}_{t}\leftarrow\textsc{AllReduce}(S_{t}^{(k)})$
12:          if $H(\overline{\mathbf{S}}_{t})>\Theta$ then
13:               $\mathbf{w}_{t}^{(k)}\leftarrow\textsc{AllReduce}(\mathbf{w}^{(k)}_{t})$triangleright In-place
Equation (4) indicates that the model variance estimation is reduced to estimating $\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$ efficiently. Computing $\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$ exactly can be done by an AllReduce operation, but this would incur a communication cost equal to performing synchronization, since the dimension of model drifts is equal to that of the models.
We next present two techniques for communication-efficient variance estimation. To present them in a uniform manner, we define the local state $\mathbf{S}_{t}^{(k)}$, and the average state $\overline{\mathbf{S}}_{t}$ as
$\mathbf{S}_{t}^{(k)}=\left(\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}\,,\,\mathbf{\sigma}_{t}^{(k)}\right)\quad,\quad\overline{\mathbf{S}}_{t}=\left(\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}\,,\,\frac{1}{K}\sum_{k=1}^{K}\mathbf{\sigma}_{t}^{(k)}\right)$
where vectors $\mathbf{\sigma}_{t}^{(k)}$ are lower-dimensional representations of $\mathbf{u}_{t}^{(k)}$.
In the two subsequent sections, we formally define $\mathbf{\sigma}_{t}^{(k)}$, and a function $H$ that is applied on the average state vector $\overline{\mathbf{S}}_{t}$. Each specification of $\mathbf{\sigma}_{t}^{(k)}$ and $H$ corresponds to a different FDA variant.
3.1. SketchFDA: Sketch-based Estimation
An optimal estimator for $\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$ can be obtained through the utilization of AMS sketches, as detailed in (Cormode and Garofalakis, 2005). An AMS sketch of a vector $\mathbf{v}\in\mathbb{R}^{d}$ is an $l\times m$ real matrix:
$\operatorname{sk}\left(\mathbf{v}\right)=\begin{bmatrix}\xi_{1}&\xi_{2}&\dots&\xi_{l}\end{bmatrix}^{\top}\in\mathbb{R}^{l\times m}\;\;,\;\;l\cdot m\ll d$
@@ 3.1. SketchFDA: Sketch-based Estimation
$\mathcal{M}_{2}\left(\operatorname{sk}(\mathbf{v})\right)=\mathrm{median}\left\{\left\|\xi_{i}\right\|_{2}^{2}\;,\;i=1,\dots,l\right\}$
@@
is an excellent estimator of $\left\|\mathbf{v}\right\|_{2}^{2}$, as with probability at least $\left(1-\delta\right)$, $\mathcal{M}_{2}\left(\operatorname{sk}\left(\mathbf{v}\right)\right)$ is within $\epsilon$-relative error of $\left\|\mathbf{v}\right\|_{2}^{2}$.
@@
In this approach, the lower-dimensional representation of the local drifts becomes $\sigma_{t}^{(k)}=\operatorname{sk}\left(\mathbf{u}_{t}^{(k)}\right)\in\mathbb{R}^{l\times m}$.
Theorem 3.1.
Define:
$S_{t}^{(k)}=\left(\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}\,,\,\operatorname{sk}\left(\mathbf{u}_{t}^{(k)}\right)\right)\in\mathbb{R}\times\mathbb{R}^{l\times m}$
@@
$H\left(\overline{\mathbf{S}}_{t}\right)=\frac{1}{K}\sum_{k}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\frac{1}{1+\epsilon}\mathcal{M}_{2}\left(\frac{1}{K}\sum_{k=1}^{K}\operatorname{sk}\left(\mathbf{u}_{t}^{(k)}\right)\right).$
@@
Then, the condition $H\large(\overline{\mathbf{S}}_{t}\large)\leq\Theta$ implies $\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)\leq\Theta$ with probability at least ($1-\delta$).
Proof.
$H\left(\overline{\mathbf{S}}_{t}\right)$ $=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\frac{1}{1+\epsilon}\mathcal{M}_{2}\left(\frac{1}{K}\sum_{i=1}^{K}\operatorname{sk}\left(\mathbf{u}_{t}^{(k)}\right)\right)$
${=}\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\frac{1}{1+\epsilon}\mathcal{M}_{2}\left(\operatorname{sk}\left(\frac{1}{K}\sum_{i=1}^{K}\mathbf{u}_{t}^{(k)}\right)\right)$
$=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\frac{1}{1+\epsilon}\mathcal{M}_{2}\left(\operatorname{sk}\left(\overline{\mathbf{u}}_{t}\right)\right)$
${\geq}\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}\quad\text{with prob. at least ($1-\delta$)}$
$=\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$
We proved that $H(\overline{\mathbf{S}}_{t})\geq\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$ with probability at least ($1-\delta$), i.e., we overestimate the model variance with probability at least ($1-\delta$), completing the proof.
3.2. LinearFDA: Linear Approximation
Although AMS sketches provide good estimates for variance, their dimension is in the several hundreds, and the communication cost of AllReduce on sketches, performed at each step, may be non-negligible. Therefore, we also examine a minimum-cost, ad-hoc estimation variant.
In this approach, $\mathbf{\sigma}_{t}^{(k)}$ is reduced to a scalar, that is, $\sigma_{t}^{(k)}=\left\langle\xi\,,\,\mathbf{u}_{t}^{(k)}\right\rangle\in\mathbb{R}$ where $\xi\in\mathbb{R}^{d}$ is any unit vector.
Theorem 3.2.
Define:
$\mathbf{S}_{t}^{(k)}=\left(\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}\,,\,\left\langle\xi\,,\,\mathbf{u}_{t}^{(k)}\right\rangle\right)\in\mathbb{R}\times\mathbb{R}\;\;,\;\;\left\|\xi\right\|_{2}=1$
$H\left(\overline{\mathbf{S}}_{t}\right)=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left\|\frac{1}{K}\sum_{i=1}^{K}\left\langle\xi\,,\,\mathbf{u}_{t}^{(k)}\right\rangle\right\|_{2}^{2}$
Then, the condition $H\large(\overline{\mathbf{S}}_{t}\large)\leq\Theta$ implies $\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)\leq\Theta$.
Proof.
$H\left(\overline{\mathbf{S}}_{t}\right)$ $=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left\|\frac{1}{K}\sum_{i=1}^{K}\left\langle\xi\,,\,\mathbf{u}_{t}^{(k)}\right\rangle\right\|_{2}^{2}$
$=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left\|\left\langle\xi\,,\,\frac{1}{K}\sum_{i=1}^{K}\mathbf{u}_{t}^{(k)}\right\rangle\right\|_{2}^{2}$
$=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left|\left\langle\xi\,,\,\overline{\mathbf{u}}_{t}\right\rangle\right|^{2}$
$\geq\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left\|\xi\right\|_{2}^{2}\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$
$=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$
$=\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$
We proved that $H(\overline{\mathbf{S}}_{t})\geq\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$, i.e., we always overestimate the model variance, completing the proof.
A random choice of $\xi$ is likely to perform poorly (force synchronization prematurely), as it is likely to be close to orthogonal to $\overline{\mathbf{u}}_{t}$. A good choice would be a vector correlated to $\overline{\mathbf{u}}_{t}$. A heuristic choice is to take the difference of the models of the last two synchronizations, assuming that the current model drift is correlated with the model drift at the last synchronization. All workers can compute this heuristic choice of $\xi$ without communication cost:
$\xi=\frac{\overline{\mathbf{w}}_{t_{0}}-\overline{\mathbf{w}}_{t_{-1}}}{\left\|\overline{\mathbf{w}}_{t_{0}}-\overline{\mathbf{w}}_{t_{-1}}\right\|_{2}}.$
3.3. Discussion
The main intuition for FDA is summarized in making the decision to synchronize dynamic, based on model variance during training. By Eq. 4, model variance is the difference between the average of squared (local) drifts $\frac{1}{K}\sum\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}$ and the squared average (global) drift $\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$. When both are low, or when both are high, synchronizing the local models is not necessary.
When contrasting SketchFDA and LinearFDA, it should first be noted that both methods overestimate the true model variance. However, SketchFDA is probabilistically quite accurate. Our findings suggest that setting $l=5$ and $m=250$ (as defined in Section 3.1) yields excellent accuracy in practice. These values have been adopted in our experiments and are thus recommended. With these settings, the size of a sketch is $4k$ bytes, significantly smaller than the size of all our models.
Between the two variants of FDA, LinearFDA requires less computational effort and bandwidth to compute and to transmit its local state (the norm of a vector plus the dot product of two vectors, rather than a sketch). On the other hand, SketchFDA provides in practice a very accurate estimator of the true model variance which is expected to lead to fewer synchronizations.
An alternative point of view of FDA, relates to the well-known “bias vs. variance” trade-off (Goodfellow et al., 2016), which encapsulates the tension between under-fitting and over-fitting. Note that the term “variance” here refers to change to the learned model $\mathbf{w}$ when the training data changes, and is different from our model variance $\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$.